OpenStudy (anonymous):

I don't really understand the problem: if f(x) is continuous on [1,3] and 2

OpenStudy (anonymous):

\[2\le f(x) \le 4 and \int\limits_{1}^{3}F(x) \]

OpenStudy (agent0smith):

\[\large 2\le f(x) \le 4 \] The way you have it doesn't look right... are you sure it's \[\large \int\limits\limits_{1}^{3}F(x) dx\] or \[\large \int\limits\limits_{1}^{3}f(x) dx\]

OpenStudy (anonymous):

The bottom one. I pushed the shift key on accident.

OpenStudy (agent0smith):

Oh, alright then. The maximum value of the integral will be when f(x) is it's maximum value. See if you can see why... which has the greatest area? remember that the integral from 1 to 3 means the area under the curve, and these are f(x)=4 and f(x)=2 https://www.google.com/search?q=y%3D4%2C+y%3D2&aq=f&oq=y%3D4%2C+y%3D2&aqs=chrome.0.57j0l3j62l2.925j0&sourceid=chrome&ie=UTF-8

OpenStudy (anonymous):

 Hope this works :-)